# Do Similar Entities have Similar Embeddings?

### v1

@@ Abstract
Knowledge graph embedding models (KGEMs) developed for link prediction learn vector representations for graph entities, known as embeddings. A common tacit assumption is the KGE entity similarity assumption, which states that these KGEMs retain the graph’s structure within their embedding space, i.e., position similar entities close to one another. This desirable property make KGEMs widely used in downstream tasks such as recommender systems or drug repurposing. Yet, the alignment of graph similarity with embedding space similarity has rarely been formally evaluated. Typically, KGEMs are assessed based on their sole link prediction capabilities, using ranked-based metrics such as Hits@K or Mean Rank. This paper challenges the prevailing assumption that entity similarity in the graph is inherently mirrored in the embedding space. Therefore, we conduct extensive experiments to measure the capability of KGEMs to cluster similar entities together, and investigate the nature of the underlying factors. Moreover, we study if different KGEMs expose a different notion of similarity. Datasets, pre-trained embeddings and code are available at: https://github.com/nicolas-hbt/similar-embeddings.
@@ 1 Introduction
We call this the KGE entity similarity assumption. Answering this question initially requires remembering that most embedding-based models are trained to maximize rank-based metrics for link prediction, which disregard any semantic considerations. One could then argue that the maximization of such metrics is at least partially decoupled from the task of learning similar vectors for similar entities. However, as related entities are more likely to appear in similar triples (e.g., featuring the same predicate), it is reasonable to believe that semantically close entities – especially those connected to other entities through a shared set of predicates – are also more likely to be assigned similar vectors [23]. This raises our second research question:
@@
In other words, we ask whether KGEMs with good link prediction performance w.r.t. rank-based metrics such as Mean Reciprocal Rank (MRR) and Hits@$K$ necessarily group similar entities close in the embedding space. If so, it would be sensible to study which of the two is most likely to influence the other, and whether a form of causality (rather than just correlation) between these two aspects exists.
@@
  • We show that different KGEMs turn their attention to different subsets of predicates for learning similar embeddings for related entities in the KG. This suggests that the notion of similarity in the embedding space is partially influenced by the distribution of predicates in the close neighborhood around KG entities.
The remainder of the paper is structured as follows. Related work about KGEMs and their use in semantic-related tasks is presented in Section 2. In Section 3, we detail our general approach for measuring similarity. Section 4 details our experimental setting. Experimental results are provided and discussed in Section 5. Lastly, Section 6 summarizes the key findings and outlines directions for future research.
@@ 2 Related Work
Analzying the KGE entity similarity assumption. While many of the approaches implicity rely on the KGE entity similarity assumption to hold, there are only few works actually validating this assumption. Portisch et al. [23] provide anecdotic results for a few embedding approaches, also suggesting that the underlying notions of similarity might differ, but do not conduct any formal evaluation. Jain et al. [14] analyze KGEMs on the basis of class assignments, showing that the original class assignments can only be reconstructed to a limited extent with classification and clustering methods, which rather questions the assumption of similar entities being close in the vector space. A similar study is conducted by Alshargi et al. [2], also concluding that “the current quality of the embeddings for ontological concepts is not in a satisfactory state”. Inspired by those findings, this paper presents an in-depth study of the KGE entity similarity assumption.
@@ 3 Approach
In this section, we detail our proposal for quantifying the KGE entity similarity assumption. While similarity in embedding spaces is usually measured using cosine similarity, there is no uniform definition of entity similarity in KGs. In Section 3.1, we elaborate on the metrics used to capture the notion of similarity between entities in the KG. Section 3.2 discusses how the aforementioned notions of similarity can be compared.
@@ 3.1 Towards a Graph-based Notion of Similarity
In this section, we detail our attempt at capturing the notion of similarity in the original KG. Unlike the notion of similarity in the embedding space, which is typically measured as the cosine similarity of two vectors,it should be noted that graph-based similarity cannot be measured in a single, uniform way. Multiple metrics and approaches can be used, and the choice between one or another largely depends on the particular aspect to be measured. In what follows, we intend to explain the rationale behind our modeling choices, and briefly mention alternatives that we ultimately discarded.
To measure the similarity between two entities in the KG, (1) we first need a general idea of how these entities are linked, not only together, but also to other entities. Another desideratum is that (2) similarity should not be overly influenced by the connectivity of an entity. Otherwise, central entities (i.e. with many in-going and out-going edges) are similar to every other entity, and entities scarcely connected to others are similar to no one. Finally, since an entity cannot be fully described by its direct, immediate neighbors, i.e., (3) indirect dependencies also need to be captured.
@@
To satisfy our first desideratum (1), we consider the subgraphs around entities. We experiment with the 1-hop only vs. 2-hop subgraph neighborhood. The latter option addresses desideratum (3), as we also consider indirect dependencies. In Fig. 1, we give a concrete example of the 1-hop and 2-hop subgraphs for the entities: Bob and Julie. [^1] Note that we consider both ingoing as well as outgoing edges to define those neighborhoods. Now that subgraphs are considered as the basis on which similarity should be assessed, we need a metric to actually evaluate this. Graph Edit Distance (GED) [28] has been used for this intent. However, GED is NP-hard and thus computationally demanding, it comes with the need for arbitrarily defining weights for vertex and edge insertion/deletion/substitution, and we experimentally found it was sensitive to subgraphs’ sizes (which is dentrimental to desideratum (2)). Instead, in this work we consider the Jaccard coefficient over two subgraphs rooted in two different entities as a measure of their pairwise similarity in the KG. In contrast to GED, using the Jaccard coefficient is a simpler approach that does not come with the aforementioned caveats. It is formally defined as:
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (1)
The Jaccard coefficient ranges from 00 to $1$, where 00 indicates no similarity (i.e., the two subgraphs being compared have no elements in common) and $1$ indicates identical sets.
@@
In our experiments, the Jaccard coefficient is used to measure the overlap in the 1-hop and 2-hop subgraphs of entities. In particular, its value denotes how similar the respective subgraphs of two entities are, which give us insight into how much two entities are related based on graph-based information. Whether it is for 1-hop or 2-hop, the central entity is replaced by a unique placeholder or token, e.g. :dummy. Triples forming the 1-hop neighborhood are extracted as is, while for the 2-hop paths the intermediate entity is ignored. Following the subgraphs shown in Fig. 1, triples in the two 1-hop subgraphs are:
@@
When considering 2-hop information, we additionally consider the following paths which are shared between Bob’s and Julie’s subgraphs (reverse edges are considered):
(:dummy, livesIn, inCountry, Germany)
(:dummy, livesIn, inRegion, Baden-Württemberg)
[…]
Considering the 1-hop subgraphs only: $J(Bob,Julie)_{1H}=0.2$. Adding information about the 2-hop neighbors: $J(Bob,Julie)_{2H}=0.5$. Note that the intermediate entity in paths of length $2$ is removed. For instance, we consider (:dummy, livesIn, inCountry, Germany) as the full path. This allows for capturing indirect dependencies, such as Both Julie and Bob live in a city in Germany, or Both Julie and Bob know someone who works for Vodafone. This approach better fits the general and commonsense idea that two entities should already be related if they are related to entities which are similar themselves according to their respecitve 1-hop subgraphs. Moreover, keeping the intermediate entity in the path would overweigh information already counted as part of the 1-hop subgraph. For any entity in the 1-hop neighborhood of a central entity, its own 1-hop neighborhood is known and fixed. Consequently, whenever there is a match between triples in the 1-hop neighborhood of two central entities being compared, there would be as many matches in their respective 2-hop subgraphs as there are relations connecting their shared 1-hop neighbors. For example, if Bob and Julie were living in the same city, e.g. Mannheim, this information would be counted multiple times – in (:dummy, livesIn, Mannheim), (:dummy, livesIn, Mannheim, inCountry, Germany), and (:dummy, livesIn, Mannheim, inRegion, Baden-Württemberg). The resulting relatedness between Bob and Julie would be overly influenced by that single fact. Similarly, if they were living in distinct cities – which is the case in our example – their relatedness score would be unduly penalized.
@@
3.2 Embedding-based vs. Graph-based: a Different Notion of Similarity?
Based on a given entity $e$, it is possible to get its closest neighbors in the KG (Section 3.1) and in the embedding space (Section LABEL:sec:embedding-similarity). We subsequently need to measure how much the two lists of closests neighbors between these two approaches actually overlap. A common measure to compare two ranked lists is Kendall’s Tau [18]. However, it suffers from two important caveats: the two lists should be of the same size and they should contain the same set of items. Considering our experimental purposes, the latter limitation is undesirable, as there is no reason why these two lists should consistently share the same set of entities. Consequently, we use the Rank-Biased Overlap (RBO) [39] as a metric to compare the similarity of two ranked lists. Unlike Kendall’s Tau which is correlation-based, RBO is intersection-based. Most importantly, RBO can handle lists containing different items. It also allows for the weighting of rankings, giving higher importance to items at the top of the lists through tweaking the persistence parameter $p$. A small value for $p$ will only consider the first few items, whereas a larger value will encompass more items. However, we experimentally found that results are very sensitive to the choice of $p$. Besides, when comparing two ranked lists, we want to consider all of their items. Consequently, we stick to the default parameter strategy $p=1$ as proposed in an open-source implementation of RBO [^2] https://github.com/changyaochen/rbo that we used in our experiments. Formally, RBO is expressed as follows.
@@ 4.1 Knowledge Graph Embedding Models
In our experiments, we use seven popular KGEMs from different families of models: geometric-based (TransE [6], TransD [15], BoxE [1]), multiplicative (RESCAL [22], DistMult [40], TuckER [4]), and convolutional-based (ConvE [8]) models. Other models have been considered, e.g. ComplEx [37] and RotatE [35]. However, these models generate complex-valued embeddings, which requires a different approach than using simple cosine similarity. We trained these KGEMs using PyKEEN [^3] https://github.com/pykeen/pykeen with the provided configuration files, when available. For those datasets with no reported best hyperparameters, we used the hyperparameters reported in the original paper (e.g. for YAGO4-19K, with best hyperparameters found in [11]) or performed manual hyperparameter search and kept the sets of hyperparameters leading to the best results on the validation sets (e.g. for AIFB). We also trained RDF2Vec [24], which is a versatile embedding approach that can be adapted for different downstream applications and that relies on path-based information to encode entities (only). Even though RDF2Vec is not specifically designed to handle the link prediction task and is not evaluated w.r.t. it, its entity embeddings are expected to reflect their similarity or relatedness in the KG to some extent [23]. We used the implementation provided by pyRDF2Vec [^4] https://github.com/IBCNServices/pyRDF2Vec with the default hyperparameters reported in [23]: embeddings of dimension $200$, with $2,000$ walks maximum, a depth of $4$, a window size of $5$, and $25$ epochs for training word2vec [21] with the continuous skip-gram architecture.
@@ 4.2 Datasets
Since we also want to analyze results on a per-class basis, we consider only benchmark datasets which also come with a schema of multiple classes and relations. We use the following datasets in our evaluation: AIFB [5], Codex-S and Codex-M [27], DBpedia50 [29], FB15K-237 [36], and YAGO4-19K [11]. Entity types are not directly available in the original repository of DBpedia50555https://github.com/bxshi/ConMask and FB15k-237666https://www.microsoft.com/en-us/download/details.aspx?id=52312. However, we ran SPARQL queries against DBpedia to get entity types for DBpedia50. For FB15K-237, we reused the entity-typed version presented in [11]. Table 1 provides finer-grained statistics for these datasets.

### v2

@@ Abstract
Knowledge graph embedding models (KGEMs) developed for link prediction learn vector representations for entities in a knowledge graph, known as embeddings. A common tacit assumption is the KGE entity similarity assumption, which states that these KGEMs retain the graph’s structure within their embedding space, i.e., position similar entities within the graph close to one another. This desirable property make KGEMs widely used in downstream tasks such as recommender systems or drug repurposing. Yet, the relation of entity similarity and similarity in the embedding space has rarely been formally evaluated. Typically, KGEMs are assessed based on their sole link prediction capabilities, using ranked-based metrics such as Hits@K or Mean Rank. This paper challenges the prevailing assumption that entity similarity in the graph is inherently mirrored in the embedding space. Therefore, we conduct extensive experiments to measure the capability of KGEMs to cluster similar entities together, and investigate the nature of the underlying factors. Moreover, we study if different KGEMs expose a different notion of similarity. Datasets, pre-trained embeddings and code are available at: https://github.com/nicolas-hbt/similar-embeddings/.
@@ 1 Introduction
We call this the KGE entity similarity assumption. Notably, there is no universally accepted definition for entity similarity. In this work, we follow a straightforward approach: two entities in a KG are similar if we make similar statements about them. This aligns with the assumption of distributional semantics: words appearing in similar contexts are semantically similar. Answering RQ1 requires remembering that most embedding-based models are trained to maximize rank-based metrics for link prediction, which disregards semantics. One could then argue that maximizing such metrics is at least partially decoupled from the task of learning similar vectors for similar entities. However, as related entities are more likely to appear in similar triples (e.g., featuring the same predicate), it is reasonable to believe that semantically close entities – especially those connected to other entities through a shared set of predicates – are also more likely to be assigned similar vectors [23]. This raises our second research question:
@@
In other words, we ask whether KGEMs with good link prediction performance w.r.t. metrics such as Mean Reciprocal Rank (MRR) and Hits@$K$ necessarily group similar entities close in the embedding space. If so, it would be sensible to study which of the two is most likely to influence the other, and whether a form of causality (rather than just correlation) between these two aspects exists.
@@
  • We show that different KGEMs turn their attention to different predicate subsets for learning similar embeddings for related entities. This suggests that the notion of similarity in the embedding space is partially influenced by the predicate distribution in the close neighborhood around KG entities.
The remainder of the paper is structured as follows. Related work about KGEMs and their use in semantic-related tasks is presented in Section 2. Section 3 elaborates on our approach for measuring similarity. Section 4 details our experimental setting. Results are provided and discussed in Section 5. Lastly, Section 6 summarizes the key findings and outlines directions for future research.
@@ 2 Related Work
Analzying the KGE entity similarity assumption. While many of the approaches implicity rely on the KGE entity similarity assumption to hold, there are only few works actually explicitly validating this assumption. Portisch et al. [23] provide anecdotic examples for a few embedding approaches, also suggesting that the underlying notions of similarity might differ, but do not conduct any formal evaluation. Jain et al. [14] analyze KGEMs on the basis of class assignments, showing that the original class assignments can only be reconstructed to a limited extent with classification and clustering methods, which rather questions the assumption of similar entities being close in the vector space. A similar study is conducted by Alshargi et al. [2], also concluding that “the current quality of the embeddings for ontological concepts is not in a satisfactory state”. While Portisch et al. provide anecdotic examples, and Jain et al. and Alshargi et al. only look at class assignments, this work is the first one to empirically study the relation of entity similarity and similarity in the KGE space. Moreover, our study is more fine-grained than the previous ones, which end at the class level: while those only inspect whether entities of the same class have similar embeddings (e.g., two movies are embedded closer than a movie and a person), we also consider similarity within a class (e.g., analyze whether similar movies are embedded closer than less similar ones).
@@ 3 Approach
In this section, we detail our proposal for quantifying the KGE entity similarity assumption. While similarity in embedding spaces is usually measured using cosine similarity, there is no uniform definition of entity similarity in KGs. In Section 3.1, we elaborate on the metrics used to capture the notion of similarity between entities in the KG. Section 3.2 discusses how the aforementioned notions of similarity can be compared. It is worth noting that the line of research closest to ours is the one of Jain et al. [14]. Both [14] and our work question common assumptions that are taken for granted in the KG community. However, [14] looks at the capability of KGEMs to learn the semantics of classes, and considers the specific tasks of entity classification and clustering. In contrast, we are concerned with similarity measures of different embedding models and how they reflect entity similarity. As such, even though our work fits within an existing literature on the concept of semantic capture in embeddings, to the best of our knowledge this paper is the first one to present an approach to thoroughly investigate the extent to which entity similarity in KGs is mirrored in the embedding space.
@@ 3.1 Towards a Graph-based Notion of Similarity
In this section, we detail our attempt at capturing the notion of similarity in the original KG. Unlike the notion of similarity in the embedding space, which is typically measured as the cosine similarity of two vectors, it should be noted that graph-based similarity cannot be measured in a single, uniform way. As previsouly said, there is no universal definition for entity similarity in KGs. Multiple metrics and approaches can be used, and the choice between one or another largely depends on the particular aspect to be measured. In what follows, we intend to explain the rationale behind our modeling choices, and briefly mention alternatives that we ultimately discarded.
To measure the similarity between two entities $e_{1}$ and $e_{2}$ in one KG, we (1) determine the set of common statements about $e_{1}$ and $e_{2}$, i.e., relations to other entities that $e_{1}$ and $e_{2}$ have in common. While many entities will have such relations in common with central entities (i.e., those with many ingoing and outgoing edges), we also (2) need to make sure that centrality does not skew our similarity measure. Finally, (3) since an entity cannot be fully described by its immediate neighbors, indirect dependencies also need to be captured.
@@
To satisfy our first desideratum (1), we compare the subgraphs around $e_{1}$ and $e_{2}$. We experiment with the 1-hop only vs. 2-hop subgraph neighborhood. The latter option addresses desideratum (3), as we also consider indirect dependencies. In Fig. 1, we give a concrete example of the 1-hop and 2-hop subgraphs for entities Bob ($e_{1}$) and Julie ($e_{2}$). [^1] We consider both ingoing and outgoing edges to define those neighborhoods. The similarity of $e_{1}$ and $e_{2}$ can now be measured by the similarity of their respective subgraphs. Graph Edit Distance (GED) [28] has been used for this intent. However, GED is NP-hard and thus computationally demanding, it comes with the need for arbitrarily defining weights for vertex and edge insertion/deletion/substitution, and we experimentally found it was sensitive to subgraphs’ sizes, which is dentrimental to desideratum (2). We considered other metrics such as Katz centrality and the common-neighbors metric. However, as the name suggests, Katz is a measure of centrality, not similarity. Besides, it takes into account all the paths between two entities and is therefore sensitive to the absolute number of paths. Common neighbors does not consider predicates and is only suited for unlabeled graphs. [^2] Given, e.g., (SAP, headquarter, Germany), (BOSCH, headquarter, Germany), (Berlin, capitalOf, Germany), with common neighbors, SAP, BOSCH, and Berlin would be equally similar.
@@
In our experiments, we use the Jaccard coefficient is used to measure the overlap in the 1-hop and 2-hop subgraphs of entities. In particular, its value denotes how similar the respective subgraphs of two entities are, which gives us insight into how much two entities are related based on graph-based information. The central entity ($e_{1}$ or $e_{2}$) is replaced by a unique token, e.g. :dummy. Triples forming the 1-hop neighborhood are extracted as is, while for the 2-hop paths the intermediate entity is ignored (written as $\Box$ below).
More formally: let $\mathcal{T}_{1}(A)$ be the set of all triples $\{(A,p,o)\cup(s,p,A)\}$, where $A$ is replaced by a dummy identifier. Then, $J_{1}(e_{1},e_{2})$ is the Jaccard overlap of $\mathcal{T}_{1}(e_{1})$ and $\mathcal{T}_{1}(e_{2})$, i.e., $\frac{|\mathcal{T}_{1}(e_{1})\cap\mathcal{T}_{1}(e_{2})|}{|\mathcal{T}_{1}(e_{1})\cup\mathcal{T}_{1}(e_{2})|}$.
Let $\mathcal{T}_{2}(A)$ be the set of all paths $\{(A,p1,\Box,p2,o)\cup(s,p1,\Box,p2,A)\}$, where $A$ is replaced by a dummy identifier. Then $J_{2}(e_{1},e_{2})$ is the Jaccard overlap of $\mathcal{T}_{2}(e_{1})$ and $\mathcal{T}_{2}(e_{2})$, i.e., $\frac{|\mathcal{T}_{2}(e_{1})\cap\mathcal{T}_{2}(e_{2})|}{|\mathcal{T}_{2}(e_{1})\cup\mathcal{T}_{2}(e_{2})|}$.
Following Fig. 1, triples in the two 1-hop subgraphs are:
@@
When considering 2-hop information, we also consider the following paths which are shared between Bob’s and Julie’s subgraphs (reverse edges included):
(:dummy, livesIn, $\Box$, inCountry, Germany)
(:dummy, livesIn, $\Box$, inRegion, Baden-Württemberg)
(:dummy, friend, $\Box$, staffMember-1, Vodafone)
The similarity of Bob and Julie considering the 1-hop subgraphs only is $J_{1}(Bob,Julie)=0.2$ (since they share one out of five triples). Considering the 2-hop neighborhood graphs: $J_{2}(Bob,Julie)=0.5$ (since they share three out of six two-hop paths). For instance, we consider (:dummy, livesIn, $\Box$, inCountry, Germany) as the full path. This allows for capturing indirect dependencies, such as both Julie and Bob live in a city in Germany, or both Julie and Bob know someone who works for Vodafone. This approach better fits the general and commonsense idea that two entities should already be similar if they share relations to entities which are similar themselves according to their respecitve 1-hop subgraphs. Moreover, keeping the intermediate entity in the path would overweigh information already counted as part of the 1-hop subgraph, especially if the intermediate entity were a high degree entity (desideratum (3)). For any entity in the 1-hop neighborhood of a central entity, its own 1-hop neighborhood is known and fixed. Consequently, whenever there is a match between triples in the 1-hop neighborhood of two central entities being compared, there would be as many matches in their respective 2-hop subgraphs as there are relations connecting their shared 1-hop neighbors. For example, if Bob and Julie were living in the same city, e.g. Mannheim, this information would be counted multiple times – in (:dummy, livesIn, Mannheim, inCountry, Germany), (:dummy, livesIn, Mannheim, inRegion, Baden-Württemberg), etc.. The resulting relatedness between Bob and Julie would thus be overly influenced by that single fact. Similarly, if they were living in distinct cities – which is the case in our example – their relatedness score would be unduly penalized.
@@
3.2 Embedding vs. Graph: a Different Notion of Similarity?
Based on a given entity $e$, it is possible to get its $N$ closest neighbors in the KG (using the $J_{1}$ and $J_{2}$ metrics defined above) and in the embedding space (using cosine similarity). To assess if the two notions of similarity are different, we need to measure how much the two lists of closests neighbors between these two approaches actually overlap.
A common measure to compare two ranked lists is Kendall’s Tau [18]. However, it suffers from two important caveats: the two lists should be of the same size and they should contain the same set of items. Considering our experimental purposes, the latter limitation is undesirable, as there is no reason why these two lists (of the top $N$ similar entities according to J1 or J2, and in the embedding space) should share the same set of entities.
Consequently, we use the Rank-Biased Overlap (RBO) [39] to compare the similarity of two ranked lists. Unlike Kendall’s Tau which is correlation-based, RBO is intersection-based. Most importantly, RBO can handle lists containing different items. It also allows for the weighting of rankings, giving higher importance to items at the top of the lists through tweaking the persistence parameter $p$. A small value for $p$ will only consider the first few items, whereas a larger value will encompass more items. However, we experimentally found that results are very sensitive to the choice of $p$. Besides, when comparing two ranked lists, we want to consider all of their items. Consequently, we stick to the default parameter strategy $p=1$ as proposed in an open-source implementation of RBO [^3] https://github.com/changyaochen/rbo/ that we used in our experiments. Formally, RBO is expressed as follows.
@@ 4.1 Knowledge Graph Embedding Models
In our experiments, we use seven popular KGEMs from different families of models: geometric-based (TransE [6], TransD [15], BoxE [1]), multiplicative (RESCAL [22], DistMult [40], TuckER [4]), and convolutional-based (ConvE [8]) models. Other models have been considered, e.g. ComplEx [37] and RotatE [35]. However, these models generate complex-valued embeddings, which requires a different approach than using simple cosine similarity. We trained these KGEMs using PyKEEN [^4] https://github.com/pykeen/pykeen/ with the provided configuration files, when available. For those datasets with no reported best hyperparameters, we used the hyperparameters reported in the original paper (e.g. for YAGO4-19K, with best hyperparameters found in [11]) or performed manual hyperparameter search and kept the sets of hyperparameters leading to the best results on the validation sets (e.g. for AIFB). We also trained RDF2Vec [24], which is a versatile embedding approach that can be adapted for different downstream applications and that relies on path-based information to encode entities (only). Even though RDF2Vec is not specifically designed to handle the link prediction task and is not evaluated w.r.t. it, its entity embeddings are expected to reflect their similarity or relatedness in the KG to some extent [23]. We used the implementation provided by pyRDF2Vec [^5] https://github.com/IBCNServices/pyRDF2Vec/ with the default hyperparameters reported in [23]: embeddings of dimension $200$, with $2,000$ walks maximum, a depth of $4$, a window size of $5$, and $25$ epochs for training word2vec [21] with the continuous skip-gram architecture.
@@ 4.2 Datasets
Since we also want to analyze results on a per-class basis, we consider only benchmark datasets which also come with a schema of multiple classes and relations. We use the following datasets in our evaluation: AIFB [5], Codex-S and Codex-M [27], DBpedia50 [29], FB15K-237 [36], and YAGO4-19K [11]. Entity types are not directly available in the original repository of DBpedia50666https://github.com/bxshi/ConMask/ and FB15k-237777https://www.microsoft.com/en-us/download/details.aspx?id=52312/. However, we ran SPARQL queries against DBpedia to get entity types for DBpedia50. The resulting class hierarchy is a subset of the DBpedia ontology, with a maximum depth of $8$, an average fan-out (branching factor) of $3.40$, and each entity being typed with an average of $4.55$ classes. For FB15K-237, we reused the entity-typed version presented in [11]. Table 1 provides finer-grained statistics for these datasets.